# A Fast and Optimal Learning-based Path Planning Method for Planetary Rovers

### v2

@@ 5.2 The Effects of Different Gaussian Positional Encoding
In contrast to other works in the field, we feed the start and end point information, encoded via positional encoding, into the model through separate channels. This approach effectively ensures that the model accords sufficient weight to the locational information of the start and end points during prediction. Literaturekulvicius2021one also inputs the start and end points via separate channels, but only with a single pixel each. While the approach can effectively predict paths in small grid maps up to ${20\times 20}$, it struggles in larger and more complex scenes such as the elevation maps discussed herein. The experiments depicted in Figure. 3 demonstrate that appropriate positional encoding of the start and end point parameters can considerably enhance the model’s delineation of traversable areas, whereas utilizing single pixels without positional encoding leads to unsatisfactory performance.
@@
In Figure. 3, ${G}$ represents ${Gaussian}$ position encoding and ${E}$ represents ${Euclidean}$ position encoding. ${1/7}$ indicates that ${\sigma}$ is taken as ${1/7}$of the map size in Eq. 5. The Euclidean encoding method calculates the Euclidean distance from each grid cell to the start or end point as the relative position encoding, which is then normalized to generate the final 2D encoding. It can be seen that the model prediction performs the best when ${\sigma}$ is around ${1/4}$ of the map size. When ${\sigma}$ is very small, it is close to no position encoding, resulting in a significant decrease in model prediction performance.
@@ 5.3 Application to Dynamic Environment
For autonomous navigation robots, global path planning and local path planning target different scenarios, which are based on the global map from external sensors and local map from onboard sensors, respectively. However, there is no fundamental difference between them. The same algorithm with proper modifications can be applied to both scenarios. The NNPP method introduced in this paper essentially learns the mapping relationship between the terrain fluctuations within a certain area measured by traversability cost and the optimal path given start and end points. Therefore, whether it is a global map or a local map, the NNPP is capable of predicting optimal path regions from uneven terrains. To validate this idea, in this section, we employ the NNPP trained on a global lunar map dataset for path planning in local Martian maps. To further simulate the presence of obstacles in the local maps, several artificial noise points are introduced into the local Martian Digital Elevation Model (DEM), as illustrated in Figure. 8.
Furthermore, the predictive performance of the model in dynamic environments is also evaluated. The red pentagrams represent the destination points, while the green rectangles represent the starting points. Figure. 8 ${a)}$ and ${e)}$ validate the path planning for reaching different destination points under the same scene and starting point, while Figure. 8 ${b)-d)}$ and ${f)-h)}$ validate the path planning for the same starting and destination points in varying scenes.
@@ 6 CONCLUSION AND FUTUREWORK
This paper presents a neural network-based approach for path planning on uneven planetary surfaces, offering a solution to the demand for fast path planning in the autonomous navigation of planetary rovers. Drawing inspiration from real-time semantic segmentation models, the study adopts a channel-wise separation of maps, starting points, and destination points specifically for the path planning scenario. Furthermore, the investigation explores the impact of different positional encoding methods on the predictive performance of the model. The final simulation experiments demonstrate that the proposed method in this paper significantly reduces the computation time required for searching the optimal path in an elevation map compared to the traditional ${A^{\star}}$ method. Without the need for GPU acceleration, the model achieves an inference speed of 17 frames per second (FPS), enabling CPU-based real-time path planning in computationally constrained vehicular systems operating in dynamic scenarios. The inference time can be observed in Figure. 5.

### v3

@@ 5.2 The Effects of Gaussian Positional Encoding
Unlike previous approaches, we input start and end point information, encoded using positional encoding, into the model via distinct channels. This approach effectively ensures that the model accords sufficient weight to the locational information of the start and end points during prediction. Literature [35] also inputs the start and end points via separate channels, but only with a single pixel each. While the approach can effectively predict paths in small grid maps up to ${20\times 20}$, it struggles in larger and more complex scenes such as the elevation maps discussed herein. The experiments depicted in Figure LABEL:fig:eps3 demonstrate that appropriate positional encoding of the start and end point parameters can considerably enhance the model’s delineation of traversable areas, whereas utilizing single pixels without positional encoding leads to unsatisfactory performance.
Figure 13: The impact of different positional encoding methods on model performance. In the horizontal axis, ${G}$ represents ${Gaussian}$ positional encoding, ${1/7}$ denotes the ${\sigma}$ value equal to one-seventh of the map size, and ${E}$ indicates ${Euclidean}$ encoding.
@@
In Figure 13, ${G}$ represents ${Gaussian}$ position encoding and ${E}$ represents ${Euclidean}$ position encoding. ${G:1/7}$ indicates that ${\sigma}$ is taken as ${1/7}$ of the map size in Eq.7. The Euclidean encoding method calculates the Euclidean distance from each grid cell to the start or end point as the relative position encoding, which is then normalized to generate the final 2D encoding. It can be seen that the model prediction performs the best when ${\sigma}$ is around ${1/4}$ of the map size. When ${\sigma}$ is very small, it is close to no position encoding, resulting in a significant decrease in model prediction performance.
@@ 5.3 Application to Dynamic Environment
For autonomous navigation robots, global path planning and local path planning target different scenarios, which are based on the global map from external sensors and local map from onboard sensors, respectively. Both global and local traversability cost maps derived from DEM data share the same format (${H\times W\times 1}$). However, they differ in two key aspects:
  • 1.
    (1) The real-world resolution corresponding to each grid cell. Figure 12 depicts a global map with a resolution of 20 meters, while the local map in Figure 14 has a resolution of 0.2 meters.
  • 2.
    (2) The presence of occlusion artifacts in local maps due to the limited field of view of onboard sensors.
@@
We argue that, the NNPP model effectively addresses the task definition challenges outlined in Section 3.3, regardless of whether the task is global or local. The NNPP method introduced in this paper essentially learns the mapping relationship between the terrain fluctuations within a certain area measured by traversability cost and the optimal path given start and end points. Therefore, whether it is a global map or a local map, the NNPP is capable of predicting optimal path regions from uneven terrains. To validate this idea, in this section, we employ the NNPP trained on a global lunar map dataset for path planning in local Martian maps. To further simulate the presence of obstacles in the local maps, several artificial noise points are introduced into the local Martian Digital Elevation Model (DEM), as illustrated in Figure 14.
Figure 14: The results of model-based path planning in dynamic scenarios are presented, including scenarios with different endpoints for the same starting point ${(a)}$ and ${(e)}$, as well as scenarios with the same starting and ending points but changing maps ${(b)-(d)}$ and ${(f)-(h)}$.
Figure 14 ${(a)}$ and ${(e)}$ validate the path planning for reaching different destination points under the same scene and starting point, while Figure 14 ${(b)-(d)}$ and ${(f)-(h)}$ validate the path planning for the same starting and destination points in varying scenes. Highlighting the adaptability of NNPP to dynamic environments due to its fast speed (17FPS), we further demonstrate its capabilities through additional experiments involving dynamic environments. Specifically, we introduce two obstacle regions into the traversability cost map, each assigned a traversal cost of 1 (hereafter referred to as obstacles).
As illustrated by the Figure 15 below, the NNPP model exhibits spatial consistency in its predictions. Despite the varying obstacle positions in timestamps $T$ and $T+4\Delta t$, the NNPP model generates identical segmentation regions. This indicates that the NNPP model can effectively distinguish whether the location of obstacles influences the optimal path. In the demonstration scenario, the upper obstacle moves to the left, and the lower obstacle moves to the right. With the start and end points remaining unchanged, the NNPP model accurately estimates the optimal path distribution for each frame and then combines it with the ${A^{\star}}$ algorithm to determine the optimal path for the current frame. This demonstrates the NNPP model’s ability to handle path planning in dynamic environments.
Figure 15: The figures illustrate the map changes from time $T$ to $T+4\Delta t$, with the white rectangular obstacles at the top and bottom representing the added areas with a traversal cost of 1. The upper obstacle moves to the left, and the lower obstacle moves to the right. The first column shows the heuristic region changes (colored in cyan) output by the NNPP model for the five time steps with fixed start and goal points. The second column shows the graphing region (colored in cyan) for path planning using the ${A^{\star}}$ algorithm under the same conditions.
${D^{\star}}$ and ${A^{\star}}$ are considered fundamental algorithms in the field of path planning, with ${D^{\star}}$ being more commonly employed in dynamic environments. We believe that integrating the heuristic region computed by the NNPP model into either the ${A^{\star}}$ or ${D^{\star}}$ algorithm can significantly enhance performance. The path planning algorithm generates a trajectory based on the current map and start and end points. The agent then navigates along this path until reaching the destination. If the map remains static, replanning is unnecessary. This scenario represents a static environment. When the map changes during path execution, causing the original path to become suboptimal or obstructed, the algorithm replans the path based on the updated map and the current start and end points. Incremental search algorithms like ${D^{\star}}$ efficiently generate new paths by utilizing information from previous iterations, adapting to dynamic environments without replanning from scratch.
Figure 16: (a) to (e) illustrate the map changes at five different time steps. At each time step, the NNPP model is invoked to recompute the heuristic region for the current start and goal positions. (f) depicts the actual path traversed from the initial start to goal, considering the dynamic changes in the environment.
In this paper, we aim to demonstrate that the NNPP model, when combined with the foundation algorithm, can also achieve significant performance improvements for dynamic path planning problems.
Table 3: Path planning results for four dynamic time steps. AT represents path planning time, CC represents cumulative traverse cost, and ↑ indicates that the larger the value, the better the performance.
Frames ${A^{\star}}$ ${A^{\star}NN}$ ${D^{\star}}$ ${D^{\star}NN}$
AT↓ CC↓ AT↓ CC↓ AT↓ CC↓ AT↓ CC↓
$T$ 1.77 20.61 0.42 20.61 1.73 20.61 0.42 20.61
$T+\Delta t$ 0.77 15.81 0.25 17.18 0.62 15.81 0.20 16.05
$T+2\Delta t$ 0.33 10.62 0.17 11.62 0.24 10.62 0.17 10.62
$T+3\Delta t$ 0.17 6.82 0.10 6.82 0.09 6.82 0.11 6.98
[…]
To illustrate the effectiveness of our approach, we conducted the following experiment. As shown in Figure 16, a path was planned from the yellow rectangular start point to the red star-shaped goal point. Assuming the agent could perfectly follow the planned path, it reached the yellow location in (b) after traversing 45 grids. At this point, the environment map changed, and the previously planned path was blocked by an obstacle. Therefore, the path from the current position to the goal was replanned. After another 45 steps, the environment changed again, and the second planned path was blocked. A third planning was performed, resulting in the path shown in (c). After 45 more steps, the environment changed again. While the previously planned path (the path in (c)) could still lead to the goal without collision, its cumulative traversal cost was not optimal. Therefore, the current environment was re-planned to obtain the optimal path shown in (d). The agent followed this path to reach the goal in (e). Figure 16 (f) shows the actual path traversed from the initial start point to the goal.
By incorporating the heuristic region of the NNPP model into the ${D^{\star}}$ algorithm’s path search at each iteration, we obtain the ${D^{\star}}$ method, which employs the same mechanism as the ${A^{\star}}$ method.
Table 4: Path planning results for five map samples are presented. The ’-s’ flag indicates a static scene, while the ’-d’ flag indicates a dynamic scene.
scenarios ${A^{\star}}$ ${A^{\star}NN}$ $D^{\star}$ ${D^{\star}NN}$
AT↓ CC↓ AT↓ CC↓ MM↑ AT↓ AT↓ CC↓
map1-s 2.37 120.51 0.85 122.87 0.05
map1-d 4.13 143.87 1.46 146.75 0.05 3.08 1.13 146.57
[…]
map2-d 3.97 161.23 1.33 164.41 0.05 2.85 1.11 164.52
[…]
map3-d 3.65 125.24 1.16 126.64 0.05 2.74 1.09 126.81
[…]
map4-d 3.72 136.82 1.24 139.55 0.07 2.89 1.12 139.55
[…]
map5-d 3.58 134.42 1.12 136.10 0.07 2.64 0.98 135.97
The experimental data corresponding to the specific timestamps in Figure 16 are summarized in Table.3 including the time consumption and accumulated cost ($CC$) values for path planning using the ${A^{\star}}$, ${A^{\star}NN}$, ${D^{\star}}$, and ${D^{\star}NN}$ algorithms. For the initial path planning, ${A^{\star}}$ and ${D^{\star}}$ algorithms exhibit similar computational costs due to their underlying principles. While ${A^{\star}}$ and ${D^{\star}}$ guarantee optimal paths, ${D^{\star}}$ leverages past information to significantly reduce computation time in subsequent planning tasks. This efficiency is further enhanced by ${D^{\star}NN}$, which achieves a roughly 2.8×2.8$\times$2.8 ×speedup compared to ${D^{\star}}$, as shown in the Table.3.
At time $T+3\Delta t$, ${D^{\star}NN}$ takes 0.11s to compute the path, while ${D^{\star}}$ algorithm takes only 0.09s. This difference arises from ${D^{\star}NN}$’s two-step computation process: first, it calls a pre-trained NNPP model to predict the heuristic region, and then it employs a foundation algorithm to plan the optimal path within the reduced area. As illustrated in Figure 9, NNPP’s inference time averages around 0.06s. Consequently, when the start and end points are close together, NNPP may fall behind foundation algorithms. However, in most task scenarios, NNPP consistently delivers significant performance improvements.
We designed dynamic scenarios similar to Figure 16 for the five map samples in Figure 12. Each map sample was calculated five times and averaged to obtain Table.4. Based on the information presented in Table.4, we can draw the following four conclusions:
  • 1.
    In static scenarios across five map samples, the ${A^{\star}NN}$/${A^{\star}}$ method averaged 0.78s/2.32s, while in dynamic scenarios, the average time for both methods was 1.26s/3.81s. Overall, NNPP achieved a 3$\times$×speedup over the ${A^{\star}}$ algorithm in both static and dynamic scenarios.
  • 2.
    In dynamic scenarios, the average time consumption of the ${D^{\star}}$ algorithm is 2.84 seconds. This demonstrates that the acceleration contribution of NNPP on ${A^{\star}}$ in dynamic scenarios even exceeds that of the ${D^{\star}}$ algorithm without NNPP acceleration.
  • 3.
    In dynamic scenarios, the average time consumption of the ${D^{\star}NN}$ method is 1.08s, which shows that NNPP can also bring about a performance improvement of nearly 2.5$\times$×speedup for the ${D^{\star}}$ algorithm.
  • 4.
    The NNPP method is expected to result in a 1%-2% decrease in path quality.
@@ 6 CONCLUSION AND FUTUREWORK
This paper proposes NNPP, a learning-based model for computing heuristic regions, which further accelerates optimal path planning on uneven terrain. Drawing inspiration from real-time semantic segmentation models, the study adopts a channel-wise separation of maps, starting points, and destination points specifically for the path planning scenario. Furthermore, the investigation explores the impact of different positional encoding methods on the predictive performance of the model. The final simulation experiments demonstrate that the proposed method in this paper significantly reduces the computation time required for searching the optimal path in an elevation map compared to the traditional ${A^{\star}}$ method. Without the need for GPU acceleration, the model achieves an inference speed of 17 FPS, enabling CPU-based real-time path planning in computationally constrained vehicular systems operating in dynamic scenarios. The inference time can be observed in Figure 9.